Metro Area New York Chicago Washington, D.C. Los Angeles San Francisco Houston Boston Dallas-Fort Worth Philadelphia Denver Seattle Minneapolis-St. Paul San Jose Source: USA Today Number of Recent College Graduates Who Moved to the Area from 2000–2011 (thousands) 200 97 92 92 64 51 51 50 49 37 34 32 27

present data on a variety of topics. For each data set described in boldface, find the a. mean. b. median. c. mode (or state that there is no mode). d. midrange.
Answer #1

(a)

Mean () is given by:

So,

Mean is:

67.3846

(b)

Arranging numbers in ascending order, we get:

27,32,34,37,49,50,51,51,64,92,92,97,200

n = 13

Median =(13+1)/2th item = 7th item = 51.

So,

Median is:

51

(c)

The modes are:

51 and 92

because 51 and 92 occur 2 times

(d)

midrange = (High + Low)/2 = (200 + 27)/2 = 113.5

So

midrange is:

113.5